# Khoroshkin–Tolstoy approach for quantum superalgebras

### v2

@@ 5. Cartan–Weyl generators
Let us discuss the difference with the case of the quantum algebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M}))$, see, for example, the paper [52]. First of all note that the expression for the universal $R$-matrix is invariant with respect to the normalization of the Cartan–Weyl generators, see section 6. We use a normalization that is convenient for calculations. The main difference from the case of $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M}))$ is related to the presence of isotropic roots. Here we use a hint following from the form of the defining relations for the generators of the second Drndfeld’s realization of $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ [20, 53].
6. Universal $R$-matrix
$\Pi(a_{1}\otimes a_{2})=(-1)^{[a_{1}][a_{2}]}a_{2}\otimes a_{1}.$
One can show that the mapping
$\Delta^{\prime}=\Pi\circ\Delta$
It should be noted that we define the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ as a $\mathbb{C}$-algebra. It can be also defined as a $\mathbb{C}[[\hbar]]$-algebra, where $\hbar$ is considered as an indeterminate. In this case there exists an element $\mathcal{R}$ of $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))\otimes\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ connecting the two comultiplications in the sense that
$\Delta^{\prime}(a)=\mathcal{R}\,\Delta(a)\mathcal{R}^{-1}$ (6.1)
see [9] for the case of a general quantum algebra, and [21] for a generalization to the case of quantum superalgebras. For the meaning of the superscripts in the above relations see, for example, appendix A.1 of the paper [52]. The element $\mathcal{R}$ is called the universal $R$-matrix. One can show that it satisfies the universal Yang-Baxter equation
$\mathcal{R}^{(12)}\,\mathcal{R}^{(13)}\,\mathcal{R}^{(23)}=\mathcal{R}^{(23)}\,\mathcal{R}^{(13)}\,\mathcal{R}^{(12)}$ (6.3)
in $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))^{\otimes 3}$.
Using two representations, $\varphi_{1}$ and $\varphi_{2}$, of the quantum algebra one construct the corresponding $R$-operator $R_{\varphi_{1}\varphi_{2}}(\zeta_{1},\zeta_{2})$ as
$R_{\varphi_{1}\varphi_{2}}(\zeta_{1},\zeta_{2})=(\varphi_{1\zeta_{1}}\otimes\varphi_{2\zeta_{2}})(\mathcal{R}).$
For any three representations, $\varphi_{1}$, $\varphi_{2}$ and $\varphi_{3}$, we get the three $R$-operators satisfying the usual operator Yang–Baxter equation
$R^{(12)}_{\varphi_{1}\varphi_{2}}(\zeta_{1},\zeta_{2})R^{(13)}_{\varphi_{1}\varphi_{3}}(\zeta_{1},\zeta_{3})R^{(23)}_{\varphi_{2}\varphi_{3}}(\zeta_{2},\zeta_{3})=R^{(23)}_{\varphi_{2}\varphi_{3}}(\zeta_{2},\zeta_{3})R^{(13)}_{\varphi_{1}\varphi_{3}}(\zeta_{1},\zeta_{3})R^{(12)}_{\varphi_{1}\varphi_{2}}(\zeta_{1},\zeta_{2}).$
For the notations used, we again refer to the paper [52].
In our case, the universal $R$-matrix exists only in some restricted sense, see the papers [54, 2, 3] for the case of quantum algebras $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M}))$, and the discussion below for the case of $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$. Here, an $R$-operator is constructed in the following way. Let $V_{1}$ and $V_{2}$ be finite dimensional $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$-modules, and $\varphi_{1}$ and $\varphi_{2}$ be the corresponding representations. Define the $R$-operator $R_{\varphi_{1}\varphi_{2}}(\zeta_{1},\zeta_{2})$ as
$R_{\varphi_{1}\varphi_{2}}(\zeta_{1},\zeta_{2})=\rho_{\varphi_{1}\varphi_{2}}(\zeta_{1}|\zeta_{2})\,(\varphi_{\zeta_{1}}\otimes\varphi_{\zeta_{2}})(\mathcal{R}_{\prec\delta}\,\mathcal{R}_{\sim\delta}\,\mathcal{R}_{\succ\delta})\,K.$ (6.4)
Here $\rho_{\varphi_{1}\varphi_{2}}(\zeta_{1}|\zeta_{2})$ is a scalar normalization factor, $\mathcal{R}_{\prec\delta}$, $\mathcal{R}_{\sim\delta}$ and $\mathcal{R}_{\succ\delta}$ are elements of the superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))\otimes\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$, while $K$ is an element of $\mathrm{End}(V_{1}\otimes V_{2})$.
where
$q_{\gamma}=(-1)^{[\gamma]}q^{(\gamma|\gamma)}.$
It is useful to have in mind that
$(-1)^{[\alpha_{ij}+n\delta]}=(-1)^{[(\delta-\alpha_{ij})+n\delta]}=d_{i}d_{j},\quad(\,(\delta-\alpha_{ij})+n\delta\,|\,(\delta-\alpha_{ij})+n\delta\,)=d_{i}+d_{j}.$
The quantities $a_{\gamma}$ are determined from the equation
$[\![e_{\gamma},\,f_{\gamma}]\!]=a_{\gamma}\frac{q^{h_{\gamma}}-q^{-h_{\gamma}}}{q-q^{-1}}.$
Here for $\gamma=\sum_{i=0}^{M+N-1}m_{i}\alpha_{i}$ we put
$h_{\gamma}=\sum_{i=0}^{M+N-1}d_{i}m_{i}h_{i}.$
In particular, $h_{\delta}=c$, therefore,
$q^{h_{\delta}}=q^{c}=1.$
The element $\mathcal{R}_{\sim\delta}$ is defined as
$\mathcal{R}_{\sim\delta}=\exp\big{(}-(q-q^{-1})\sum_{n\in\mathbb{Z}_{>0}}\,\sum_{i,j=1}^{M+N-1}(-1)^{n}o_{i}^{n}o_{j}^{n}d_{i}d_{j}U_{nij}\,(e_{n\delta,\,\alpha_{i}}\otimes f_{n\delta,\,\alpha_{j}})\big{)},$
$[\![e_{\alpha_{i}+m\delta},\,e_{n\delta;\alpha_{j}}]\!]=d_{j}\,o_{j}^{n}T_{nij}\,e_{\alpha_{i}+(m+n)\delta},$ (6.5)
and
$o_{i}=(-1)^{i-1},\quad i<M,\qquad o_{i}=(-1)^{i},\quad i\geq M.$
The endomorphism $K$ is defined as follow. Let $v_{1},v_{2}\in V$ be weight vectors of weights $\lambda_{1}$ and $\lambda_{2}$ respectively. Then we have
$K(v_{1}\otimes v_{2})=q^{-\sum_{i,\,j=1}^{l}d_{i}d_{j}C_{ij}\,\langle\lambda_{1},\,h_{i}\rangle\langle\lambda_{2},\,h_{j}\rangle}\,(v_{1}\otimes v_{2}),$ (6.6)
where $C_{ij}$ are the matrix elements of the matrix $C$ inverse to the symmetrized Cartan matrix $B$ of the Lie superalgebra algebra $\mathfrak{sl}_{M|N}$. For the explicit form of $C_{ij}$ we again refer to the appendix B.
One can show that it is possible to work with the $R$-operators $R_{\varphi_{1}\varphi_{2}}(\zeta_{1},\zeta_{2})$ defined by (6.4) as if it were defined by the universal $R$-matrix satisfying equations (6.1) and (6.2). In particular, they satisfy the operator Yang–Baxter equation.
@@ 7. Evaluation modules
In applications to the theory of quantum integrable systems, one usually considers families of representations parametrized by a complex parameter called a spectral parameter. We introduce a spectral parameter in the following way. Assume that the quantum loop algebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ is $\mathbb{Z}$-graded. It means that $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ expands into a direct sum
@@
$\Gamma_{\zeta_{1}\zeta_{2}}=\Gamma_{\zeta_{1}}\circ\Gamma_{\zeta_{2}}$ (7.1)
for any $\zeta_{1},\zeta_{2}\in\mathbb{C}^{\times}$. Now, for any representation $\varphi$ of $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ we define the corresponding family $\varphi_{\zeta}$ of representations as
@@
If $V$ is the $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$-module corresponding to the representation $\varphi$, we denote by $V_{\zeta}$ the $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$-module corresponding to the representation $\varphi_{\zeta}$.
The common way to endow $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ by a $\mathbb{Z}$-gradation is to assume that
@@
where $s_{i}$ are arbitrary integers. We denote
$s=\sum_{i=0}^{M+N-1}s_{i},$ (7.2)
and assume that $s$ is non-zero. It is clear that for such a $\mathbb{Z}$-gradation one has
@@
To construct representations of $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$, it is common to use the Jimbo’s homomorphism $\epsilon$ from the quantum loop superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ to the quantum superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ defined by the equations
$\epsilon(q^{\nu h_{0}})=q^{-\nu(K_{M+N}+K_{1})},$ $\epsilon(q^{\nu h_{i}})=q^{\nu(K_{i}-d_{i}d_{i+1}K_{i+1})},$ (7.4)
$\epsilon(e_{0})=-F_{1,\,M+N}\,q^{d_{1}K_{1}+d_{M+N}K_{M+N}},$ $\epsilon(e_{i})=E_{i,\,i+1},$ (7.5)
$\epsilon(f_{0})=q^{-d_{1}K_{1}-d_{M+N}K_{M+N}}E_{1,\,M+N},$ $\epsilon(f_{i})=F_{i,\,i+1},$ (7.6)
where $i$ runs from $1$ to $M+N-1$. It is clear that if $\pi$ is a representation of $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ then $\pi\circ\epsilon$ is a representation of the quantum loop algebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$. We call such representations and the corresponding modules evaluation ones. Starting with the representation $\pi$ of $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ defined by equations (3.10) and (3.11), we define a family of evaluation representations
@@
of $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$. Using explicit form of the Jimbo’s homomorphism, we obtain
@@
$\varphi_{\zeta}(f_{0})=\zeta^{-s_{0}}q^{-1}\mathbb{E}_{1,\,M+N},$ $\varphi_{\zeta}(f_{i})=\zeta^{-s_{i}}\mathbb{E}_{i+1,\,i},$ $i=1,\ldots,M+N-1.$

### v3

@@ 5. Cartan–Weyl generators
Let us discuss the difference with the case of the quantum algebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M}))$, see, for example, the paper [53]. First of all note that the expression for the universal $R$-matrix is invariant with respect to the normalization of the Cartan–Weyl generators, see section 7. We use a normalization that is convenient for calculations. The main difference with the case of $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M}))$ is related to the presence of isotropic roots. Here we use a hint following from the form of the defining relations for the generators of the second Drindfeld’s realization of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ [20, 54].
@@ 6. Evaluation modules
In applications to the theory of quantum integrable systems, one usually considers families of representations parameterized by a complex parameter called the spectral parameter. We introduce the spectral parameter in the following way. We assume that the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ is $\mathbb{Z}$-graded. It means that $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ expands into a direct sum
@@
$\Gamma_{\zeta_{1}\zeta_{2}}=\Gamma_{\zeta_{1}}\circ\Gamma_{\zeta_{2}}$
for any $\zeta_{1},\zeta_{2}\in\mathbb{C}^{\times}$. Now, for any representation $\varphi$ of $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$, we define the corresponding family $\varphi_{\zeta}$ of representations as
@@
The common way to endow $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ with a $\mathbb{Z}$-grading is to assume that
@@
where $s_{i}$ are arbitrary integers. It is clear that for such a $\mathbb{Z}$-grading we have
@@
To construct representations of $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$, it is common to use the Jimbo’s homomorphism $\epsilon$ from the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ to the quantum superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ defined by the equations
$\epsilon(q^{\nu h_{0}})=q^{-\nu(-d_{M+N}K_{M+N}+d_{1}K_{1})},$ $\epsilon(q^{\nu h_{i}})=q^{\nu(K_{i}-d_{i}d_{i+1}K_{i+1})},$
$\epsilon(e_{0})=-F_{1,\,M+N}\,q^{d_{1}K_{1}+d_{M+N}K_{M+N}},$ $\epsilon(e_{i})=E_{i,\,i+1},$
$\epsilon(f_{0})=q^{-d_{1}K_{1}-d_{M+N}K_{M+N}}E_{1,\,M+N},$ $\epsilon(f_{i})=F_{i,\,i+1},$
where $i$ runs from $1$ to $M+N-1$. It is clear that if $\pi$ is a representation of the quantum superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ then $\pi\circ\epsilon$ is a representation of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$. We call such representations and the corresponding modules evaluation ones. Starting with the representation $\pi$ the quantum superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ defined by equations (3.1) and (3.2), we define a family of evaluation representations
@@
of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$. Using the explicit form of the Jimbo’s homomorphism, we obtain
@@
$\varphi_{\zeta}(f_{0})=\zeta^{-s_{0}}q^{-1}\mathbb{E}_{1,\,M+N},$ $\varphi_{\zeta}(f_{i})=\zeta^{-s_{i}}\mathbb{E}_{i+1,\,i},$ $i=1,\ldots,M+N-1.$ (6.5)
By definition, the representations $\varphi_{\zeta}$ act on the superspace $\mathbb{C}_{M|N}$.
Let $V$ be a $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$-module. The generators $q^{x}$, $x\in\widetilde{\mathfrak{h}}$, form an abelian group in $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$. Let a vector $v\in V$ be a common eigenvector for all operators $\varphi(q^{x})$, then
$q^{x}v=q^{\langle\mu,\,x\rangle}v$
for some unique element $\mu\in\widetilde{\mathfrak{h}}^{*}$. Using the first relation of (4.2), we obtain
$q^{\nu c}v=q^{\nu\langle\mu,\,c\rangle}v=v$
for any $\nu\in\mathbb{C}$. Therefore, the element $\mu$ satisfies the equation
$\langle\mu,\,c\rangle=0,$
and there is a unique element $\lambda\in\mathfrak{h}^{*}$ such that $\mu=\widetilde{\lambda}$. For the definition of $\widetilde{\lambda}$ see section  4. This leads to the following definition. A $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$-module $V$ is said to be a weight module if
$V=\bigoplus_{\lambda\in\mathfrak{h}^{*}}V_{\lambda},$
where
$V_{\lambda}=\{v\in V\mid q^{x}v=q^{\langle\widetilde{\lambda},\,x\rangle}v\mbox{ for any }x\in\widetilde{\mathfrak{h}}\}.$
This means that any vector of $V$ has the form
$v=\sum_{\lambda\in\mathfrak{h}^{*}}v_{\lambda},$
where $v_{\lambda}\in V_{\lambda}$ for any $\lambda\in\mathfrak{h}^{*}$, and $v_{\lambda}=0$ for all but finitely many of $\lambda$. The space $V_{\lambda}$ is called the weight space of weight $\lambda$, and a nonzero element of $V_{\lambda}$ is called a weight vector of weight $\lambda$. One says that $\lambda\in\mathfrak{h}^{*}$ is a weight of $V$ if $V_{\lambda}\neq\{0\}$. Any finite dimensional $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$-module is a weight module.
For any $i=1,\ldots,M+N-1$ the weight $\omega_{i}$ defined by the equation
$\langle\omega_{i},\,h_{j}\rangle=\delta_{ij}$
is called fundamental. One can show that
$\omega_{i}=\sum_{k=1}^{M+N-1}\alpha_{k}\,c_{ki}\,d_{i},$ (6.6)
where $c_{ki}$ are matrix entries of the matrix $C$ inverse to $B$.
It is not difficult to verify that for the $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$-module defined by equations (6.2)–(6.5) the basis vector $v_{i}\in\mathbb{C}_{M|N}$ is a weight vector of weight
$\lambda_{i}=\omega_{1}-\sum_{k=1}^{i-1}\alpha_{k}.$ (6.7)
7. Universal $R$-matrix
$\Pi(a_{1}\otimes a_{2})=(-1)^{[a_{1}][a_{2}]}a_{2}\otimes a_{1}.$
One can show that the mapping
$\Delta^{\prime}=\Pi\circ\Delta$
It should be noted that we define the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ as a $\mathbb{C}$-algebra. It can be also defined as a $\mathbb{C}[[\hbar]]$-algebra, where $\hbar$ is considered as an indeterminate. In this case there exists an element $\mathcal{R}\in\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))^{\otimes 2}$ connecting the comultiplications in the sense that
$\Delta^{\prime}(a)=\mathcal{R}\,\Delta(a)\mathcal{R}^{-1}$ (7.1)
see [9] for the case of a general quantum algebra, and [21] for the generalization to the case of quantum superalgebras. For the meaning of the superscripts in the above relations see, for example, appendix A.1 of the paper [53]. The element $\mathcal{R}$ is called the universal $R$-matrix. One can show that it satisfies the universal Yang-Baxter equation
$\mathcal{R}^{(12)}\,\mathcal{R}^{(13)}\,\mathcal{R}^{(23)}=\mathcal{R}^{(23)}\,\mathcal{R}^{(13)}\,\mathcal{R}^{(12)}$
in $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))^{\otimes 3}$.
Using a representation $\varphi$ of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$, one construct the corresponding $R$-operator $R_{\varphi}(\zeta_{1},\zeta_{2})$ as
$R_{\varphi}(\zeta_{1},\zeta_{2})=(\varphi_{\zeta_{1}}\otimes\varphi_{\zeta_{2}})(\mathcal{R}).$
It satisfies the usual operator Yang–Baxter equation
$R^{(12)}_{\varphi}(\zeta_{1},\zeta_{2})R^{(13)}_{\varphi}(\zeta_{1},\zeta_{3})R^{(23)}_{\varphi}(\zeta_{2},\zeta_{3})=R^{(23)}_{\varphi}(\zeta_{2},\zeta_{3})R^{(13)}_{\varphi}(\zeta_{1},\zeta_{3})R^{(12)}_{\varphi}(\zeta_{1},\zeta_{2}).$ (7.3)
For the notations used, we again refer to the paper [53].
In our case, the universal $R$-matrix exists only in some restricted sense, see the papers [55, 2, 3] for the case of quantum algebras $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M}))$, and the discussion below for the case of $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$. Here, an $R$-operator is constructed in the following way. Let $V$ be a finite dimensional $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$-modules, and $\varphi$ be the corresponding representation. Define the $R$-operator $R_{\varphi}(\zeta_{1},\zeta_{2})$ as
$R_{\varphi}(\zeta_{1},\zeta_{2})=\rho_{\varphi}(\zeta_{1},\zeta_{2})\,(\varphi_{\zeta_{1}}\otimes\varphi_{\zeta_{2}})(\mathcal{R}_{\prec\delta}\,\mathcal{R}_{\sim\delta}\,\mathcal{R}_{\succ\delta})\,K_{\varphi}.$ (7.4)
Here $\rho_{\varphi}(\zeta_{1},\zeta_{2})$ is a scalar normalization factor, $\mathcal{R}_{\prec\delta}$, $\mathcal{R}_{\sim\delta}$ and $\mathcal{R}_{\succ\delta}$ are elements of the superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))^{\otimes 2}$, while $K_{\varphi}$ is an element of the algebra $\mathrm{End}(V^{\otimes 2}V)$.
where
$q_{\gamma}=(-1)^{[\gamma]}q^{(\gamma|\gamma)}.$
It is useful to have in mind that
$[\alpha_{ij}+n\delta]=[(\delta-\alpha_{ij})+n\delta]=[i]+[j],\quad(\,(\delta-\alpha_{ij})+n\delta\,|\,(\delta-\alpha_{ij})+n\delta\,)=d_{i}+d_{j}.$
The quantities $a_{\gamma}$ are determined from the equation
$[\![e_{\gamma},\,f_{\gamma}]\!]=a_{\gamma}\frac{q^{h_{\gamma}}-q^{-h_{\gamma}}}{q-q^{-1}},$
where for $\gamma=\sum_{i=0}^{M+N-1}m_{i}\alpha_{i}$ we put
$h_{\gamma}=\sum_{i=0}^{M+N-1}d_{i}m_{i}h_{i}.$
In particular, $h_{\delta}=c$, therefore,
$q^{h_{\delta}}=q^{c}=1.$
The element $\mathcal{R}_{\sim\delta}$ is defined as [^7] In fact, in our case $[d_{i}]_{q}=d_{i}$. We write $[d_{i}]_{q}$ to present expressions valid for a more general case.
$\mathcal{R}_{\sim\delta}=\exp\big{(}-(q-q^{-1})\sum_{n\in\mathbb{Z}_{>0}}\,\sum_{i,j=1}^{M+N-1}(-1)^{n}u_{nij}[d_{i}]_{q}[d_{j}]_{q}(e_{n\delta,\,\alpha_{i}}\otimes f_{n\delta,\,\alpha_{j}})\big{)},$
$[\![e_{\alpha_{i}+m\delta},\,e_{n\delta;\,\alpha_{j}}]\!]=[d_{j}]_{q}^{-1}\,o_{i}^{n}o_{j}^{n}t_{nij}\,e_{\alpha_{i}+(m+n)\delta}.$ (7.5)
The quantities $o_{i}$ have the form
$o_{i}=(-1)^{i-1},\quad i<M,\qquad o_{i}=(-1)^{i},\quad i\geq M.$
In contrast to the case of ordinary quantum algebras, the general form of the quantities $o_{i}$ for quantum superalgebras is unknown. We found them using matching considerations by selection.
The operator $K_{\varphi}$ is defined as follow. Let $v,w\in V$ be weight vectors of weights $\lambda,\mu\in\mathfrak{h}^{*}$, respectively. Then we have
$K_{\varphi}(v\otimes w)=q^{-(\,\lambda\,|\,\mu\,)}v\otimes w.$ (7.6)
One can show that it is possible to work with the $R$-operator $R_{\varphi}(\zeta_{1},\zeta_{2})$ defined by equation (7.4) as if it were defined by the universal $R$-matrix, satisfying equations (7.1) and (7.2). In particular, it satisfies the operator Yang–Baxter equation (7.3).